Let X be a continuous random variable with density function below.

"x = \begin{cases}
 \frac{x}{2}, &\text{for } 0 ≤ x ≤ 2 \\
 0, &\text{otherwise } 
\end{cases}"

Find E [|X-E[X]|].


1
Expert's answer
2021-06-30T06:43:09-0400

Solution:

"f(x) = \begin{cases}
 \frac{x}{2}, &\text{for } 0 ≤ x ≤ 2 \\
 0, &\text{otherwise } 
\end{cases}"

"E[X]=\int_0^2x(\dfrac x2)dx=\int_0^2(\dfrac {x^2}2)dx
\\=[\dfrac{x^3}6]_0^2
\\=\dfrac 86-0
\\=\dfrac 43"

Now, "E[|X-E[X]|]=E[|X-\dfrac 43|]"

"=\int_0^2|x-\dfrac 43|(\dfrac x2) dx
\\=\int_0^{\frac 43}(\dfrac 43-x)(\dfrac x2) dx+\int_{\frac 43}^2(x-\dfrac 43)(\dfrac x2) dx"

"=\int_0^{\frac 43}(\dfrac 23x-\dfrac {x^2}2) dx+\int_{\frac 43}^2(-\dfrac 23x+\dfrac {x^2}2) dx"

"=[(\dfrac {x^2}3-\dfrac {x^3}6)]_0^{\frac 43}+[(-\dfrac {x^2}3+\dfrac {x^3}6)]_{\frac 43}^2"

"=[(\dfrac {({\frac 43})^2}3-\dfrac {({\frac 43})^3}6)-0]+[(-\dfrac {2^2}3+\dfrac {2^3}6)]-[(-\dfrac {({\frac 43})^2}3+\dfrac {({\frac 43})^3}6)]"

"=[(\dfrac {{\frac {16}9}}3-\dfrac {{\frac {64}{27}}}6)]+[-\dfrac {4}3+\dfrac {8}6]-[-\dfrac {{\frac {16}9}}3+\dfrac {{\frac {64}{27}}}6]"

"=[\dfrac {16}{27}-\dfrac {32}{81}]+[0]-[-\dfrac {16}{27}+\dfrac {32}{81}]
\\=2[\dfrac {16}{27}-\dfrac {32}{81}]
\\=\dfrac {32}{81}"